Laura hit a volleyball into the air with an initial upward velocity of 24 feet per second. The height h in feet of the ball above the ground can be moduled by
h=16t^2 24t+3, where t is the time in seconds above the ground.Find the time it takes the ball to reach 12 feet above the ground.

Answer by stanbon(75887)

h= -16t^2 + 24t+3, where t is the time in seconds above the ground.Find the time it takes the ball to reach 12 feet above the ground.
-----
The "h" is height; so let h = 12 and solve for "t":
------
-16t^2+24t+3 = 12
----
16t^2-24t+9 = 0
(4t-3)(4t-3) = 0
---
t = 0.75 seconds
==========================
Cheers,
Stan H.